geerky42 (geerky42):

\[\Huge \int \sqrt{4x^2+1} ~ dx\]

geerky42 (geerky42):

I don't know where to start...

OpenStudy (zarkon):

trig sub

geerky42 (geerky42):

With what?

OpenStudy (anonymous):

\[x={1\over2}\tan\theta\]

OpenStudy (anonymous):

or \[ x={1\over2}\sinh t\\ dx={1\over2}\cosh t dt\\ I=\int \sqrt{\sinh^2t+1}{1\over2}\cosh tdt\\ I={1\over2}\int \cosh^2t dt\\ I={1\over2}\int {\cosh2t+1\over2}dt\\ \vdots \]